Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

  

Sales (13,000 units × $20 per unit) $ 260,000
Variable expenses 130,000
Contribution margin 130,000
Fixed expenses 145,000
Net operating loss $ (15,000 )

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,700 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $82,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $39,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.50 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,900?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $53,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,300 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,300)?
Answer #1

Answer 1

CM Ratio 50.00%
Break-even point in unit sales       14,500
Break-even point in dollar sales $ 290,000

Calculations:

CM Ratio =130,000/260,000
Break-even point in unit sales =145,000/10
Break-even point in dollar sales =145,000/50%

Answer 2

Increase by $ 34,300

Calculations:

Increase by =(82,000*50%) - 6,700

Answer 3

Profit $ 24,000.00

Calculations:

Before After
Sales =13000*20 =26000*18
Less: Variable cost =13000*10 =26000*10
Less: Fixed cost 145000 =145000+39000

Answer 4

Unit sales to attain target profit    15,778.95

Calculations:

Unit sales to attain target profit =(4,900+145,000)/9.5

Answer 5

CM Ratio 65.00%
Break-even point in unit sales       15,230.77
Break-even point in dollar sales $ 304,615.38

Total:

Sales $   406,000.00 $   406,000.00
Less: Variable cost $   203,000.00 $   142,100.00
Contribution $   203,000.00 $   263,900.00
Less: Fixed cost $   145,000.00 $   198,000.00
Profit $     58,000.00 $     65,900.00

Per Unit:

Sales $            20.00 $            20.00
Less: Variable cost $            10.00 $               7.00
Contribution $            10.00 $            13.00
Less: Fixed cost $               7.14 $               9.75
Profit $               2.86 $               3.25

Yes, the company should automate its operations.

Calculations:

CM Ratio =13/20
Break-even point in unit sales =198,000/13
Break-even point in dollar sales =198,000/65%

in case of any doubt, please comment.